# Can Uncertainty Quantification Enable Better Learning-based Index Tuning?

### v1

@@ 5.3 Results on Uncertainty Quantification
Second, the quality of uncertainty quantification is related to the base model and dataset. Across three benchmarks, the uncertainty modeling quality of the AMA base model consistently surpasses that of the LIB model, particularly evident in the proposed AE and Hybrid models. This may be attributed to the fact that the model’s ability to fit the data also determines its accuracy and capability in UQ. As illustrated in Table V, AMA’s accuracy is significantly higher than LIB’s. Accurate results necessitate effective encoding of inputs into a hidden vector, the foundational principle of our methodology, which accounts for these observations.
@@
Overall, these 16 models demonstrate superior performance on the TPC-H+ bencamark compared to TPC-DS+, with the poorest performance observed on the JOB+ benchmark, where the most effective UQ methods filter only 61% of Uncertainty Samples.
Third, the settings for uncertainty thresholds significantly impact the ability to detect Uncertainty Samples. Uncertainty distribution within a dataset is usually confined to a narrow range, and the relationship between the threshold and this range dictates the model’s detection capabilities. For example, in the AMA-Ensemble model under TPC-H+, setting a filtering threshold at the 81th percentile of the training dataset excludes only 38.73% of Uncertainty Samples, whereas a threshold at the 75th percentile increases exclusion to 73.34%.
Fourth, HSTO, specifically designed for Self-Attention, did not demonstrate improved performance on this task. Neither model achieved optimal performance across the three benchmarks.
Fifth, among the three MCD-based methods, MCD-all outperforms the others, achieving optimal results in five out of six scenarios. MCD-dpp provides a richer sampling diversity compared to MCD-last and outperforms it in the majority of scenarios.
TABLE IV: Time for model inference per batch (ms).
@@
LIB 2.76 2.30 5.58 AMA 0.18 0.18 0.25
BNN 61.85 44.46 50.75 BNN 10.77 7.41 10.54
MCD-all 184.59 45.03 155.14 MCD-all 6.19 2.94 4.90
MCD-last 60.00 89.63 133.04 MCD-last 5.23 4.13 5.23
MCD-dpp 60.01 147.86 83.63 MCD-dpp 260.82 78.39 51.45
Ensemble 22.72 14.15 29.90 Ensemble 1.82 0.80 1.71
AE 3.08 4.76 6.37 AE 0.52 0.51 0.51
Hybrid 60.74 103.95 98.16 Hybrid 10.28 9.82 9.26
STO 151.40 137.33 157.01
HSTO 64.46 109.25 327.03
TABLE V: The impact of uncertainty quantification methods on the prediction errors. Models that have the most negative impact are marked in red, while those with the most positive impact are marked in green.
Dataset LIB BNN MCD-all MCD-last MCD-dpp Ensemble STO HSTO AE Hybrid AMA BNN MCD-all MCD-last MCD-dpp Ensemble AE Hybrid
TPC-H+ 50th 0.13 +0.10 +0.01 +0.02 +0.04 +0.00 +0.03 +0.00 +0.09 +0.05 0.01 +0.02 +0.00 +0.00 +0.02 +0.02 +0.00 +0.00
95th 0.47 +0.02 +0.00 +0.00 -0.04 +0.00 +0.00 +0.00 +0.05 +0.01 0.05 +0.06 +0.01 +0.03 +0.00 +0.03 +0.00 +0.02
99th 0.65 -0.11 -0.05 -0.04 -0.07 -0.01 +0.00 -0.05 -0.12 -0.01 0.11 +0.07 +0.00 +0.00 +0.03 +0.01 -0.03 +0.02
TPC-DS+ 50th 0.06 +0.01 +0.00 -0.01 +0.00 +0.00 +0.00 +0.00 +0.00 +0.00 0.01 +0.01 +0.00 +0.00 +0.00 +0.00 +0.00 +0.00
95th 0.44 +0.03 +0.00 -0.02 +0.01 +0.00 +0.00 -0.02 +0.00 -0.04 0.20 +0.06 +0.02 +0.00 +0.09 +0.03 +0.01 +0.01
99th 0.71 +0.00 +0.00 +0.03 +0.00 +0.00 +0.00 +0.00 +0.02 +0.00 0.34 +0.08 +0.06 +0.05 +0.07 +0.04 +0.05 +0.07
JOB+ 50th 0.19 +0.13 +0.01 +0.03 +0.04 +0.00 +0.04 +0.05 +0.03 +0.01 0.03 +0.02 +0.00 +0.00 +0.05 +0.00 +0.00 +0.01
95th 0.57 +0.04 +0.00 +0.00 +0.00 +0.00 +0.02 +0.00 +0.03 +0.00 0.19 +0.06 -0.01 +0.00 +0.05 +0.00 +0.00 +0.04
99th 0.76 +0.16 -0.02 +0.02 +0.00 -0.02 +0.07 +0.09 +0.04 +0.04 0.37 +0.17 +0.00 +0.00 +0.11 +0.00 +0.08 +0.10
5.4 Results on Prediction Accuracy and Inference Efficiency
Table IV shows the time taken by different models to process a single batch, which includes predicting the relative error $\hat{rb}_{predicted}$ and estimating uncertainty, in milliseconds. Models in which the inference time with UQ is on the same order of magnitude as the base model are marked in green.
The primary factors influencing inference efficiency are the base model structure, model hyperparameters, and methods for UQ. LIB, utilizing heavier Multi-head Attention, requires significantly more inference time, up to an order of magnitude higher than AMA.For the models based on the same BE model, the optimal hyperparameters vary across different datasets, which can also lead to several times difference in inference efficiency. Methods like BNN, MCD, and HSTO necessitate multiple inferences on the same model, resulting in inference times that are proportionally related to the number of inferences, significantly exceeding the base model and AE by one to two orders of magnitude. Specifically, MCD-dpp introduces extra computational overhead due to sampling processes, notably affecting the faster base model AMA, and to a lesser extent, the heavier LIB model where it is less of a determining efficiency factor. The inference time for Ensemble models scales with the number of models involved; selecting five models during hyperparameter tuning has resulted in higher overall efficiency compared to other multiple-inference UQ methods. Of all methods evaluated, AE proves the most efficient, necessitating only an additional inference for the decoder network, which keeps its inference times close to those of the base model on all six benchmarks.
Table V evaluates the impact of incorporating UQ methods into BE models. We assess performance using the absolute difference between predicted and actual values, i.e. $|rb_{actual}-\hat{rb}_{predicted}|$. The table shows the 50th, 95th, and 99th percentiles of error distributions for LIB and AMA. For models to which UQ methods have been added, the table shows the change in error relative to the base model, where models that have the most negative impact are marked in red, and those with the most positive impact are marked in green.
Among all UQ methods, only BNN significantly affected the accuracy of the BE model, while the other methods had no substantial impact. Each model was retrained using 30 distinct sets of hyperparameters, so any minor variations in accuracy can be attributed to differences in hyperparameter settings. In the case of BNN, however, each neuron is replaced by a value sampled from a distribution, which effectively doubles the number of network parameters (as each deterministic neuron corresponds to values such as mean, variance, etc., in BNN) and significantly expands the hyperparameter space (e.g., requiring a prior distribution). This makes the training process fundamentally different and more complex, requiring more intensive hyperparameter tuning.
Comparing to the MCD-last, MCD-dpp had a higher effect on accuracy when AMA was used as the base model, more inference burden while more UQ capability, making it better suited for integration with heavier base models. , which might be related to the changes it made in the sampling process during inference. However, more research is needed to provide a definitive conclusion, which is beyond the scope of this study.
Among the other methods, since the AE method did not alter the original structure and parameters of the BE model after training, it theoretically should not affect the accuracy of the BE model.
6 Empirical Study on Index Tuning
6.1 Experimental Setup
In this section, we explore whether modeling uncertainty can enhance index tuning results. Therefore, we integrated the aforementioned BE model into the Beauty framework and performed index tuning on six benchmarks under different budget constraints.
6.1.1 Implementation
@@
Leis et al. [38] show that no single index configuration enumeration algorithm delivers optimal performance in all scenarios. In our experiments, we used DTA [14], which balances runtime and solution quality well across various scenarios. This algorithm has a set maximum runtime and stops after enumerating all configurations or reaching this limit. However, following Leis et al.’s [38] implementation, the algorithm completes the current seed’s configuration even if it exceeds the time limit, occasionally surpassing the maximum runtime. We adjust the maximum runtime based on workload size, as shown in Table II.
Based on the performance of various models mentioned in Section 5, we selected nine models in our experiment: LIB, LIB-Ensemble, LIB-AE, LIB-Hybrid, AMA, AMA-Ensemble, AMA-AE, AMA-Hybrid, and what-if. The thresholds for each model were set as described in Subsection 5.3. When the uncertainty value of the benefits estimated by the BE-UQ model for a sample exceeds the threshold, we use what-if tools to estimate the benefits of that sample.
6.1.2 Evaluation Metrics
We use specific metrics to assess the performance of index tuning. Given a workload $W$, $\mathcal{I}_{0}$ represents the initial set of indexes, and $\mathcal{I}$ the set of selected indexes. We measure the quality of replacing $\mathcal{I}_{0}$ with $\mathcal{I}$ by the relative improvement in $W$’s cost: $\max\left(0,\frac{B(W,\mathcal{I}_{0},\mathcal{I})}{c(W,\mathcal{I}_{0})}\right)=\max\left(0,1-\frac{c(W,\mathcal{I})}{c(W,\mathcal{I}_{0})}\right).$
On some occasions, if the selected indexes increase the cost, they must be discarded, and thus the improvement in cost is zero.
Figure 11: Cost improvements achieved by different benefit estimation models.
@@ 6.1.2 Evaluation Metrics
TABLE VI: Rankings of cost improvements achieved by models.
@@
What-if 8 8 11 1
LIB 0 0 1 10
LIB-Ensemble 1 4 4 3
LIB-AE 3 1 1 4
LIB-Hybrid 3 7 0 4
AMA 3 3 3 6
AMA-Ensemble 4 2 6 7
AMA-AE 10 7 5 0
AMA-Hybrid 4 4 5 0
6.2 Results on Index Tuning

### v2

@@ V-C Performance of Uncertainty Quantification
(3) The quality of UQ is fundamentally related to both the architecture of the base model and the benchmark. Across all three benchmarks, the UQ quality of the AMA base model consistently surpasses that of the LIB model, which is particularly evident in the proposed AE and Hybrid models. This phenomenon can potentially be attributed to the fact that a model’s capacity to effectively fit the data inherently influences both its accuracy and UQ capability. As demonstrated in Table IV, AMA’s accuracy significantly exceeds that of LIB. Accurate predictions necessitate effective encoding of inputs into a hidden vector representation, which constitutes the foundational principle of our methodology and accounts for these observed performance differentials.
@@
In summary, these $16$ models demonstrate superior performance on the TPC-H+ benchmark compared to TPC-DS+, with the least effective performance observed on the JOB+ benchmark, where even the most effective UQ methods filter only approximately $60\%$ of unreliable samples.
TABLE IV: Percentiles of BE errors produced by various BE-UQ models when using both BE-UQ models and what-if tools for BE.The best-performing models are marked in green.
Dataset LIB BNN MCD-all MCD-last MCD-dpp Ensemble STO HSTO AE Hybrid AMA BNN MCD-all MCD-last MCD-dpp Ensemble AE Hybrid What-if
TPC-H+ 50th 0.41 0.56 0.12 0.19 0.36 0.09 0.34 0.18 0.09 0.07 0.28 0.31 0.33 0.32 0.08 0.36 0.05 0.04 0.05
95th 0.92 0.77 0.97 0.97 0.94 0.77 0.68 0.91 0.68 0.59 0.74 0.72 0.73 0.76 0.68 0.73 0.53 0.42 0.60
99th 0.98 1.03 0.82 0.97 0.95 0.89 0.81 0.95 0.83 0.79 0.86 0.78 0.80 0.90 0.78 1.07 0.62 0.62 0.78
TPC-DS+ 50th 0.34 0.29 0.20 0.26 0.28 0.18 0.25 0.26 0.24 0.15 0.20 0.10 0.10 0.22 0.19 0.11 0.11 0.12 0.16
95th 0.93 0.92 0.94 0.97 0.94 0.93 0.95 0.94 0.85 0.94 0.97 0.96 0.95 0.95 0.96 0.95 0.88 0.83 0.91
99th 1.14 1.16 1.40 1.14 1.28 1.42 0.95 1.22 1.09 1.12 1.16 1.19 1.39 1.13 1.33 1.44 1.11 1.04 1.62
JOB+ 50th 0.31 0.39 0.27 0.33 0.33 0.26 0.29 0.31 0.21 0.20 0.20 0.14 0.16 0.17 0.22 0.19 0.10 0.13 0.11
95th 0.84 0.69 0.89 0.91 0.94 0.77 0.87 0.83 0.88 0.79 0.95 0.71 0.83 0.70 0.75 0.71 0.73 0.68 0.82
99th 2.07 2.08 2.02 2.21 2.14 2.08 2.09 2.19 2.05 2.17 2.52 2.07 2.25 2.23 1.91 1.94 1.78 1.85 2.20
TABLE V: Model inference time per batch (ms).
@@
LIB 2.76 6.00 5.58 AMA 0.18 0.54 0.25
BNN 61.85 156.73 50.75 BNN 10.77 20.01 10.54
MCD-all 184.59 96.80 155.14 MCD-all 6.19 12.09 4.90
MCD-last 60.00 94.72 133.04 MCD-last 5.23 11.68 5.23
MCD-dpp 60.01 121.35 83.63 MCD-dpp 260.82 183.33 51.45
Ensemble 22.72 32.48 29.90 Ensemble 1.82 3.14 1.71
AE 3.57 5.11 6.37 AE 0.52 0.87 0.51
Hybrid 60.74 81.32 98.16 Hybrid 10.28 16.45 9.26
STO 151.40 185.76 157.01
HSTO 64.46 167.11 327.03
TABLE VI: Normalized benefit estimation time of various BE-UQ models. The estimation time using only what-if tools is normalized to 100% for each benchmark.
Model TPC-H+ TPC-DS+ JOB+ Model TPC-H+ TPC-DS+ JOB+
LIB 1.51% 0.72% 0.45% AMA 100.16% 100.07% 100.01%
BNN 42.87% 36.99% 73.40% BNN 109.77% 102.41% 100.46%
MCD-all 118.38% 34.77% 17.96% MCD-all 105.56% 101.45% 100.21%
MCD-last 48.16% 33.41% 12.39% MCD-last 104.74% 101.40% 100.23%
MCD-dpp 41.51% 34.52% 10.80% MCD-dpp 336.81% 122.10% 102.27%
Ensemble 28.39% 30.91% 14.72% Ensemble 101.65% 100.37% 100.07%
AE 27.77% 23.00% 3.68% AE 100.05% 100.10% 100.02%
Hybrid 53.30% 33.70% 23.54% Hybrid 109.33% 101.98% 100.41%
STO 98.86% 47.49% 21.15%
HSTO 50.71% 43.04% 27.78%
V-D Performance of Benefit Estimation
In this subsection, we evaluate the benefit estimation (BE) performance of various BE-UQ models.
Evaluation Metric. Given a sample $(q,\mathcal{I})$ as input, the BE error produced by a BE-UQ model is
$\frac{|\hat{B}(q,\mathcal{I}_{0},\mathcal{I})-B(q,\mathcal{I}_{0},\mathcal{I})|}{c(q,\mathcal{I}_{0})},$
where $\hat{B}(q,\mathcal{I}_{0},\mathcal{I})$ is the predicted benefit, $B(q,\mathcal{I}_{0},\mathcal{I})$ is the actual benefit, and $c(q,\mathcal{I}_{0})$ is the current cost of $q$. We evaluate the BE performance of a BE-UQ model by the 50th, 95th and 99th percentiles of the BE errors produced by the BE-UQ model for all samples in $D_{eval}$. We do not use the average BE error as the evaluation metric because a few very large errors do not substantially affect the average error but significantly degrade the performance of index tuning.
Experimental Results. Table IV presents the percentiles of BE errors generated by various BE methodologies, encompassing what-if tools, learning-based BE models, and BE-UQ models. BE-UQ models augment their predictions with results from what-if tools when unreliable samples are identified, as illustrated in Figure 6. This experiment directly reflects accuracy when BE-UQ models are integrated into the index tuning process. For each benchmark, the best-performing paradigm is highlighted in green. Models derived from different base models are evaluated independently. In conjunction with Table III, two principal findings emerge:
(1) Precise uncertainty quantification substantially enhances the accuracy of benefit estimation. As demonstrated in subsection V-C, the proposed Hybrid and AE methods exhibited superior UQ performance. Consequently, in Table IV, the Hybrid approach outperforms learning-based BE models in $89\%$ of metrics, exceeds what-if tools in $72\%$ of metrics, and surpasses all other BE-UQ methods in $61\%$ of metrics. Models employing the AE technique demonstrate marginally lower accuracy than the Hybrid approach but consistently outperform other UQ methods.
(2) Deficient UQ capabilities in BE-UQ architectures combine the limitations of both learning-based BE models and what-if tools, thereby compromising the accuracy of benefit estimation. A notable example is MCD-last, which, as indicated in Table III, filters out merely $1.67\%$ of unreliable results during its worst performance. This leads to a disproportionate elimination of accurate estimation results during benefit estimation, while inaccurate results provided by the what-if tools are retained, culminating in suboptimal performance.
Our analyses demonstrate that the impact of UQ on BE accuracy is fundamentally dependent upon the robustness of UQ capabilities. When UQ capabilities are robust, BE-UQ models effectively synthesize the strengths of learning-based models and what-if tools, yielding superior results compared to either approach in isolation. Conversely, when UQ capabilities are deficient, BE-UQ accuracy typically falls between these two methods, and in certain scenarios, may underperform both approaches.
V-E Performance of Inference Efficiency
In this subsection, we evaluate the inference efficiency of various BE-UQ models.
Experimental Results (Using BE-UQ Models Only). We evaluate computational efficiency of various models by measuring the time to process a batch of samples, encompassing both benefit estimation and uncertainty quantification phases. The experimental results are presented in Table V. The BE-UQ models that achieve efficiency comparable to the base models within the same order of magnitude are highlighted in green.
(1) It is evident that among all UQ methods, AE demonstrates the highest efficiency, followed by Ensemble. Compared to the base model, AE necessitates only one additional inference for the decoder network. In contrast, the inference time for Ensemble increases linearly with the number of models. Nevertheless, Ensemble methods still demonstrate higher efficiency compared to other multiple-inference UQ techniques, such as BNN, MCD, HSTO, and Hybrid.
(2) Model inference efficiency is influenced by the base model structure and hyperparameters. For instance, LIB, which uses a heavier Multi-head Attention mechanism, requires up to an order of magnitude more inference time than AMA. For models based on the same base architecture, the optimal hyperparameters differ across various benchmarks, leading to notable differences in inference efficiency.
Experimental Results (Using Both BE-UQ Models and What-if Tools). We evaluated various models under the Beauty framework for BE with an equivalent number of samples. We recorded the BE call sequence during an index tuning process and measured the time required by what-if tools and other models to complete the BE sequence, as illustrated in Table VI. For each benchmark, we established what-if tools as the baseline, setting their estimation time to $100\%$. This assessment reflects the practical efficiency of the models when applied to index tuning tasks. The BE-UQ model with the shortest inference time is highlighted in green.
From experimental results, it is concluded that, after using the Beauty framework, the primary factor influencing BE efficiency is not the model’s inference time, but rather the number of what-if calls. This is due to the substantially lower efficiency of the what-if tools compared to learning-based models. Specifically, each what-if call requires tens to hundreds of milliseconds [1], and unlike GPU-accelerated batch processing in learning-based models, the what-if tools execute sequentially without hardware acceleration.
(1) For LIB-based models, only unreliable results from models require supplementation by what-if tools, which leads to a lower inference time compared to the what-if tools baseline. The AE remains the most efficient, but its performance improves from being one to two orders of magnitude faster than other methods to being several times faster.
(2) In contrast, models based on AMA necessitate query plans generated by what-if tools as input to obtain BE results. Consequently, when compared to baselines that utilize what-if tools, AMA-based models demonstrate marginally slower BE times in index tuning scenarios. Upon implementation of the Beauty framework, BE-UQ models incur additional computational overhead relative to baseline, primarily comprising network inference latency and UQ processing time. With the exception of MCD-dpp, which requires diversity sampling, the computational cost of the UQ for other AMA-based models remains negligible.
V-F Section Conclusion
Synthesizing the findings presented in Tables III through VI, we can draw the following conclusions: With respect to UQ capabilities, the Hybrid demonstrates superior performance, followed by the AE, both of which surpass alternative methods. This enhanced UQ capability translates into significant improvements in BE accuracy when integrated into the Beauty framework. In terms of efficiency, AE emerges as the frontrunner, outperforming both the Hybrid and other methods.
Consequently, the optimal choice between Hybrid and AE is contingent upon the specific application scenario. In scenarios where time-constrained tuning is paramount [35], AE represents the preferred option. For all other scenarios, the Hybrid approach is recommended as the primary selection.
VI Empirical Study on Index Tuning
In this section, we compare the performance of what-if tools, learning-based BE models, and BE-UQ models in index tuning, to investigate whether UQ can enhance the results of index tuning.
VI-A Experimental Setup
VI-A1 Workloads
In addition to TPC-H+, TPC-DS+, and JOB+, we also conduct index tuning on workloads containing only template-generated queries without synthetic queries, namely TPC-H, TPC-DS, and JOB.
VI-A2 Implementation
@@
Leis et al. [32] demonstrate that no single index configuration enumeration algorithm delivers optimal performance across all scenarios. In our experiments, we employed DTA [13], which effectively balances runtime and solution quality across various contexts. This algorithm operates with a predefined maximum runtime and terminates after enumerating all configurations or reaching this time constraint. However, following Leis et al.’s [32] implementation, the algorithm completes the current seed’s configuration even when exceeding the time limit, occasionally surpassing the maximum allocated runtime. We calibrate the maximum runtime according to workload size, as detailed in Table II.
Figure 9: Cost improvements achieved by different estimators.
@@
TABLE VII: Rankings of cost improvements achieved by different benefit estimators.
@@
LIB 1 1 0 21
LIB-Ensemble 2 1 5 5
LIB-AE 2 1 5 2
LIB-Hybrid 2 4 8 3
AMA 4 2 3 20
AMA-Ensemble 3 7 1 13
AMA-AE 8 8 8 7
AMA-Hybrid 15 9 6 0
What-if 5 9 8 2
TABLE VIII: Statistical information on cost improvement of different benefit estimators across various benchmarks.
Method TPC-H TPC-DS JOB TPC-H+ TPC-DS+ JOB+
Mean Var Mean Var Mean Var Mean Var Mean Var Mean Var
LIB 0.029 $7.9\cdot 10^{-4}$ 0.079 0.011 0.352 0.093 $9.5\cdot 10^{-4}$ $7.0\cdot 10^{-5}$ 0.040 0.013 0.370 0.020
LIB-Ensemble 0.041 0.002 0.216 0.012 0.555 0.019 0.001 $1.0\cdot 10^{-5}$ 0.194 0.009 0.586 0.009
LIB-AE 0.033 $3.0\cdot 10^{-4}$ 0.242 0.004 0.670 0.008 0.004 $2.4\cdot 10^{-5}$ 0.266 0.001 0.437 0.027
LIB-Hybrid 0.082 0.006 0.163 0.020 0.535 0.061 0.039 0.002 0.222 0.003 0.554 0.012
AMA 0.009 $3.3\cdot 10^{-4}$ 0.153 0.029 0.535 0.116 0.017 0.001 0.120 0.017 0.226 0.084
AMA-Ensemble 0.011 $2.4\cdot 10^{-4}$ 0.255 0.012 0.586 0.109 $0.009$ $1.6\cdot 10^{-4}$ 0.164 0.018 0.575 0.008
AMA-AE 0.085 0.005 0.258 0.011 0.542 0.185 0.059 0.002 0.287 0.002 0.602 0.002
AMA-Hybrid 0.105 0.006 0.306 0.004 0.688 0.042 0.075 0.003 0.310 0.009 0.542 0.013
What-if 0.083 0.004 0.250 0.011 0.687 0.046 0.049 $2.8\cdot 10^{-4}$ 0.251 0.009 0.493 0.078
Based on the performance of various models discussed in Section V, we selected nine models for this experiment: LIB, LIB-Ensemble, LIB-AE, LIB-Hybrid, AMA, AMA-Ensemble, AMA-AE, AMA-Hybrid, and what-if. The thresholds for each model were established as described in Subsection V-B3. When the uncertainty value of benefits estimated by the BE-UQ model for a sample exceeds the threshold, we utilize what-if tools to estimate the benefits for that particular sample.
VI-A3 Evaluation Metrics
Given a workload $W$, $\mathcal{I}_{0}$ represents the initial set of indexes, and $\mathcal{I}$ denotes the set of selected indexes. We quantify the quality of replacing $\mathcal{I}_{0}$ with $\mathcal{I}$ by the relative improvement in $W$’s cost: $\max\left(0,\frac{B(W,\mathcal{I}_{0},\mathcal{I})}{c(W,\mathcal{I}_{0})}\right)=\max\left(0,1-\frac{c(W,\mathcal{I})}{c(W,\mathcal{I}_{0})}\right).$
In scenarios where the selected index configuration leads to an increase in the total workload cost, such configurations must be discarded. Therefore, in such cases, the improvement in Figure 9 will be marked as zero. If multiple estimators have zero improvement under the same budget, the “# Worst” in Table VII will count them repeatedly.
VI-B Results on Index Tuning